Gato Inc. sells one product that sells for $52.06 and has variable costs of $37.92. The annual fixed costs are $232889 and the firm’s tax rate is 31.89%. How many units must be sold if Gato wants to earn $105869 after taxes?

Select one:

a. 23957

b. 27463

c. 16228

d. 16470
Answer #1

Revenue to be earned

= 105869/(100-31.89)

= 155438

Units needed to be sold = (Fixed cost + Revenue to be earned) /Contribution margin per unit

= (232889+155438)+(52.06-37.62)

= 388327/14.14

= 27463 (Option B)